OpenStudy (anonymous):

Determine the stopping distances for a car with an initial speed of 95 km/h and human reaction time of 1.0 s for the following accelerations.

OpenStudy (anonymous):

\[(348/abs(a))meter+26.38meter=\Delta x\]

OpenStudy (anonymous):

since the reaction time is 1s, there wont be any change in velocity during this second distance travelled i this second = initial speed * time initial speed= 95 km/h = 95 * 5 / 18= 26.388 m/s after that time the acceleration begins. so u can use the equation v^2 - u^2 = 2*a*s in this equation u consider v= 0 u= 26.388 m/s a = value of acceleration then u will get the value of s then the total stopping distance will be = (distance travelled in reaction time) + s so u can substitute the values i will do it if u give me the values of acceleration

OpenStudy (anonymous):

am i right?